OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −2 0 8 −8 2

OpenStudy (allank):

I believe the question is asking you to plug in x=-1 and y=-2 into the expression 4x-2y That will give 4(-1)-2(-2) = -4+4 = 0